OpenStudy (anonymous):

Does anyone know how to do exponential equations? 5^2x + 5^(x + 1) - 14 = 0 I am at a lose.

OpenStudy (anonymous):

I would personally take logs so you have: \[5^{2x}+5^{x+1}=14\] then take logs on both sides

OpenStudy (anonymous):

or you substitute 5^x = y then factorise and take logs

OpenStudy (anonymous):

I was using my notes and came up with 2x * ln5 + (x + 1)* ln5 = 14; but now I am stuck and do not know how to move forward from here.

OpenStudy (anonymous):

okay that is the right thing to do, but there's an easier way :) so if you let 5^x = y what would your equation become? remember \[5^{2x}+5^{x+1}-14=(5^{x})^{2}+5^{x}5^{1}-14=0\]

OpenStudy (anonymous):

Oh but don't forget with what you were doing, if you take logs, you have to do it to both sides, so the RHS would be ln(14)

OpenStudy (anonymous):

I have never seen it done this way. I think I hate pre-calculus. There seems to be a dozen ways to do the same problem.

OpenStudy (anonymous):

So if I stick with the original way that I was working the problem do I divide ln10 on both sides. I have 2x^2 +1 * ln10 = ln14.

OpenStudy (anonymous):

Right okay, let me try and work with the method that you started with (you will have to be patient as it's a long time since I did logs!) okay so we have: \[5^{2x}+5^{x+1}=14\] take logs\[(2x)\log(5)+(x+1)\log(5)=\log(14) \] \[2xlog(5) + xlog(5) + \log(5) =\log(14)\] \[3xlog(5)+\log(5)=\log(14)\] \[3xlog(5)=\log(14)-\log(5)\] \[x=\frac{ \log(14)-\log(5) }{ 3\log(5) }=\frac{ \log(14) }{ 3\log(5) }-\frac{ 1 }{ 3 }\] Do you follow that?

OpenStudy (anonymous):

Though someone else may have a clearer method, as I say, it's a while since I've used the laws of logs

OpenStudy (anonymous):

This looks like what I have been trying to do. Does log and ln mean the same thing?

OpenStudy (anonymous):

not quite, log can be to any base, where as ln is the natural logarithm, so has base 'e'

OpenStudy (anonymous):

So for a decimal approximation I got .54657. Is this anywhere near what the answer should be?

OpenStudy (anonymous):

Okay I got ~0.21325

OpenStudy (anonymous):

you need to subtract the third

OpenStudy (anonymous):

Thank you. I am not sure I have this all down correctly but at least I feel better about what I do have. This is so confusing.

OpenStudy (jack1):

\[\large 5^{2x} + 5^{x + 1} - 14 = 0 \] now add 14 to both sides \[\large 5^{2x} + 5^{x + 1} = 14\] now take log base 5 of both sides \[\large \log_{5}( 5^{2x} )+ \log_{5}(5^{x + 1}) = \log_{5}(14) \] now simplify \[\large \log_{5}(5) \times 2x+ \log_{5}(5) \times x + 1 = \log_{5}(14) \] clean up \[\large 2x+ x + 1 = \frac {\log_{5}(14)}{\log_{5}(5)} \] \[\large 3x+ 1 = \frac {\log_{5}(14)}{\log_{5}(5)} \] \[\large 3x + 1 = 1.639738 \] \[\large 3x = 0.639738 \] \[\large x = 0.213246 \] i got this... but i think im wrong...?

OpenStudy (anonymous):

Hey Jack how did you get the figures for 3x + 1 and 3x? I can not seem to come up with those same numbers.

OpenStudy (anonymous):

It doesn't need to be in base 5, it can be in any base.

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

Are you allowed to use a calculator?

OpenStudy (anonymous):

Yes.

Directrix (directrix):

5^(2x) + 5^(x + 1) - 14 = 0 Factor (5^x + 7)* (5^x -2) = 0 Zero Product Property (5^x + 7) = 0 or (5^x -2) = 0 5^x = -7 or 5^x = 2 No solution so 5^x = 2 5^x = 2 log (5^x) = log(2) x * log 5 = log 2 x = (log 2)/(log 5) x = .43.068 approx

Directrix (directrix):

Error @Jack1 Same as this: 9 + 16 = 25 3 + 4 = 5 which is not true (after taking square roots of each number in (9 + 16 = 25)

Directrix (directrix):

@Karen188 I get that you did not want to go the exponential equation route but I can't seem to make any other way work at the moment. Sorry.

OpenStudy (jack1):

ooo..."illegal" 's a bit of a strong definition there, but understanding now, cheers @Directrix

Directrix (directrix):

Didn't intend to offend - just couldn't think of a euphemism. I have made that error so many times that I think of it as illegal.

OpenStudy (jack1):

nah dude, sçool hey ;D

OpenStudy (anonymous):

Sorry guys but we all got this wrong. Thank you for all of your help. I am going to go take a nap. been up all night studying for this test and it does not look good.

Directrix (directrix):

@Karen188 After you rest, come back later and post the correct solution in this thread so that we will know how to improve our work. Thanks.

Directrix (directrix):

I already see one error in my work: x = .43.068 approx ERROR Alert Correction: x = .43068 approx And, WolframAlpha verifies: http://tinyurl.com/o9lp4a7

OpenStudy (anonymous):

I do not know the correct way to work this problem but the answers are log5(2) and .431. I hope this helps.

Directrix (directrix):

This --> log5(2) is one of the exact forms of the approximation .43068. Thanks for replying. I hope the test went well. @Karen188